OpenStudy (anonymous):

8x+2y=-4. solve for y

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

I think its 0 but im not too sure

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

isolate y like the other problems. add 8x to both sides: -4+8x = 2y. Can you go from here?

OpenStudy (anonymous):

8x+2y=-4 2y= -4 -8x y= (-4-8x)/2

OpenStudy (anonymous):

@pcnm0430 the 8 should be negative :S

OpenStudy (anonymous):

@sernst22 got it..??

OpenStudy (anonymous):

got it. thanks

OpenStudy (anonymous):

no problem.. :)